6. using the property F{f1 + f2} = F{fi} F{12} Calculate the Fourier transform of a triangular pulse. See Prob. 11. Determine the convolution of the following pulse h(t) with itself. That is, h(t) y(t) = h(t) * h(t) =?

Number one posted as a reference. Only need problem 6 answered. Thanks
Answer #1

Given ୧୪ property FCF* f2) = F(fi) x F (fa) Fouried tdarsfoom of triangular pulse. As we know the convolution of two rectangutt To I - TO A.Adu - gło to (849)“) = Y A - A du, olt<Q To t-to .:(fog)(t) = 12 (4+2 to) , -210<tco A² (2To-t) oct<9TO Iti 79Given rectangulas signal a hez) van - TE A= : convolution of he) with it self gives (hen) = p m (tf ), ctco (8-4), Octavia loๆ lolo - convolution of two rectangulas signall give triangulay signal (het) & h(t)) = triangulay signal from given propestyfourier transform of (triangular) = fale Forectangular)) ? IF( triangulay. Pulfe) = (F(2) 2 of we know that fourier thanx forBut given rectangulay signal : fourig transform of real triangulas signal 92 presiones To somente U forurier transform of tri